A boat travels 5 mph in still water. In the time it takes to travel 42 miles downstream, it can travel 12 miles upstream. Find the speed of the river.
I am having some trouble on this, but If I can find an equation or form that i could use then i think that i could solve it. Please let me know asap, I need to finish this by tomorrow. Thank You.

Answer by stanbon(75887)

A boat travels 5 mph in still water. In the time it takes to travel 42 miles downstream, it can travel 12 miles upstream. Find the speed of the river.
------------------
Downstream DATA:
distance = 42 miles ; rate = 5+r mph ; time = 42/(r+5) hrs.
------
Upstream DATA:
distance = 12 miles; rate = 5-r mph ; time = 12/(5-r) hrs.
=================
Equation:
time = time
42/(5+r) = 12/(5-r)
220-42r = 60+12r
54r = 160
r = 2.96 mph (speed of the river)
=====================================
Cheers,
Stan H.